# Achieving 70 Gb/s Over A VCSEL-Based Optical Wireless Link Using A Multi-Mode Fiber-Coupled Receiver

### v1

@@ Abstract
Index Terms:
Next-generation light fidelity (LiFi), laser-based optical wireless communication (OWC), ultra-high speed, vertical cavity surface emitting laser (VCSEL), fiber-optic photoreceiver.I Introduction
@@
Achieving data rates of tens of Gb/s using laser-based OWC for indoor applications has been actively investigated [4, 5, 6]. In [4], Lee et al. experimentally validated a dual‑wavelength wavelength division multiplexing (WDM) OWC system using a novel white‑light transmitter by integrating a nm gallium nitride (GaN) blue laser, an nm infrared (IR) laser, and a phosphor converter within a single surface‑mount device (SMD) package. The authors reported a combined data rate exceeding Gb/s over a m link at a target bit error ratio (BER) of adopted from long term evolution (LTE) systems, using two Silicon positive-intrinsic-negative (PIN) photoreceivers with a dB bandwidth of GHz to detect the optical signals of the blue and IR laser diodes. In [5], Chen et al. implemented a ‑channel LiFi link using the same SMD laser technology based on a transmitter setup that integrates ten high‑brightness, high‑bandwidth LDs, including three GaN blue LDs operating at nm, nm, and nm, and seven IR LDs with wavelegths of nm to nm. To detect the optical signals, GHz Silicon PIN photoreceivers were used, except for the nm LD, for which a GHz indium gallium arsenide (InGaAs) photoreceiver was used due to its higher responsivity. The authors demonstrated aggregate data rates in excess of Gb/s over a few meter link distance by considering a BER threshold of .
Among various types of LDs, vertical cavity surface emitting lasers are well-suited for realizing laser-based OWC due to their outstanding features. They are low-cost and low-power devices (e.g., with a few mW power consumption), and have an extensive modulation bandwidth (e.g., GHz), well-controlled output beam properties, and the ability to be fabricated as two-dimensional (2D) arrays [7]. In particular, single-mode (SM) VCSELs emit a Gaussian laser beam with a narrow spectral linewidth [8]. In addition, SM VCSELs offer a lower far field (FF) divergence angle compared to their multi-mode (MM) counterparts, providing a smaller beam spot size which leads to improved light coupling efficiency when the beam is focused on the small active area of a high speed photodiode (PD). This is of course an advantage for high speed point-to-point optical wireless link design. In [6], Osahon et al. experimentally demonstrated a line‑of‑sight optical wireless link based on an nm SM VCSEL with a -dB bandwidth of GHz, achieving a gross data rate of Gb/s over m with an eye‑safe launch power of dBm (i.e., mW), assuming a forward error correction (FEC) limit for a pre-FEC BER of . This performance was achieved by using a GHz Silicon PIN PD and spreading the data across a large number of discrete multitone (DMT) subcarriers by means of adaptive bit and power loading.
@@ II-A End-to-End System Setup
Fig. 1 shows the link configuration including the optical and optomechanical assemblies of the transmitter and receiver. We use an SM VCSEL operating at a center wavelength of nm with a narrow linewidth. This VCSEL was designed, manufactured and packaged by Vertically Integrated Systems (VIS) GmbH and was tested at the wafer level before being fully characterized for optical communications [10]. It is a multi-aperture VCSEL with a compact mini-array structure, comprising four VCSELs that are arranged in a array configuration with a small pitch of nm111Such a miniature structure is fabricated using oxidation through holes [10]. [10]. These four VCSELs are connected to a common pair of contacts such that they act as a single VCSEL. As a result, the output power is effectively increased while maintaining the SM emission spectrum [10]. The VCSEL mini-array has a nearly Gaussian beam profile with a FF full-width at half-maximum (FWHM) divergence angle of , a maximum optical power of mW, and a -dB modulation bandwidth of GHz. Furthermore, it provides a linear input dynamic range between its forward threshold current of mA and roll-over current of mA, which is suitable for DCO-OFDM modulation as discussed in the following.
@@
Fig. 2 depicts the complete experimental setup. In this work, considering the low-pass and frequency-selective characteristics of the VCSEL-based OWC system, we use DCO-OFDM with adaptive subcarrier bit and power allocation to ensure high spectral efficiency. The modulation and demodulation of the OFDM signal are implemented based on -point inverse fast Fourier transform (IFFT) and FFT operations, respectively, where , using offline digital signal processing (DSP) in MATLAB running on a laptop computer. At the transmitter, the information bits are mapped to -ary quadrature amplitude modulation (QAM) symbols to be loaded across a total of data-carrying subcarriers, with the constellation size reaching up to , depending on the available signal-to-noise ratio (SNR). The DCO-OFDM frame is constructed by concatenating the block of the complex QAM symbols with its conjugated version according to a Hermitian symmetry. A cyclic prefix (CP) of length is then added and an IFFT is performed on the resulting frame to generate the time domain OFDM signal. This signal is further shaped using a root-raised cosine filter with a roll-off factor of 0.1. The pulse-shaped digital waveform is converted into a wideband analog signal via the Keysight M8195A arbitrary waveform generator (AWG), operating at a sampling frequency of GSa/s. According to the Nyquist-Shannon theorem, the bandwidth of the signal is related to the sampling frequency via . In practice, the signal bandwidth at the AWG output can be controlled using the number of samples per symbol (SPS) parameter, denoted by , based on . We choose to obtain GHz. The AWG output signal is combined with the required direct current (DC) bias using a Mini Circuits ZX85-12G-S+ bias tee (BT) to drive the VCSEL. This BT has a GHz bandwidth with a minimum operating frequency of kHz, and an insertion loss (IL) of dB. The wideband OFDM signal is fed into the BT using a radio frequency (RF) coaxial cable with a bandwidth of GHz, a length of cm, and dB over the desired bandwidth. To collimate the light beam at the transmitter, a Thorlabs ACL5040U-B aspheric condenser lens is placed in front of the VCSEL. This lens has a diameter of mm with a clear aperture (CA) of mm, a focal length of mm, a numerical aperture (NA) of , and an anti-reflective coating (ARC) for – nm.
@@
At the receiver side, a Thorlabs C80FC-B achromatic fiber collimator is employed to collect the incident light beam from the transmitter located at a cm distance. This collimator has a CA of mm, an effective focal length of mm, an ARC for – nm matching the coating of the transmitter lens, and a Ferrule Connector/Physical Contact (FC/PC) fiber-optic output with a maximum fiber NA of . As shown in Fig. 1, the fiber collimator is mounted on a three-axis kinematic mount with high-resolution angular adjusters along with an coordinates translation stage, providing six degrees of freedom required for precise alignment of the focused beam spot on the fiber port to ensure high coupling efficiency. The collimator output is connected to the optical input of the MM fiber-coupled receiver using an OM4 graded-index fiber patch cable of length m with a core diameter of µm and FC/PC connectors. We use a Thorlabs RXM25BF fiber-optic photoreceiver, which is equipped with an InGaAs PIN PD followed by integrated transimpedance amplifier (TIA) and variable gain amplifier (VGA) stages. This receiver has a broadband responsivity in – nm, and a photodetection bandwidth of up to GHz with a minimum operating frequency of kHz. The built-in VGA helps with adjusting the transimpedance gain and the peak of the receiver frequency response to optimize the performance. There is also a digital screen on the photoreceiver to monitor the average detected photocurrent, which helps with performing the alignment using an unmodulated laser beam. The output analog signal from the photoreceiver is converted to a digital signal by the Keysight UXR0104B oscilloscope at a sampling frequency of GSa/s. The remaining DSP is performed in MATLAB. This includes frame synchronization, FFT, single-tap equalization, and QAM demapping. For each experiment, a total of OFDM frames consisting of pilot frames and data frames are sent through the OWC system to measure the BER performance.
II-B Achievable Data Rate
@@
Config. I | V | mA | µW |
Config. II | V | mA | µW |
@@ III Measurement Results
We present the measurement results for the performance of the VCSEL-based OWC system in terms of the received SNR, average BER and link data rate. The channel transfer function is estimated by sending pilot -QAM symbols of unit variance on all data-carrying subcarriers [5]. The estimated channel information is then used to determine the optimal average transmit power and constellation order of QAM symbols for each subcarrier based on the corresponding subcarrier SNR, thus adapting the transmission rate and maximizing the system spectral efficiency. In this work, the Hughes-Hartogs (HH) bit and power loading algorithm is applied [13], which is widely used in multi-carrier communication systems. The optimal DC operating point of the VCSEL is obtained through extensive measurements and two configurations are identified. These configurations along with the resulting average received optical powers, denoted by , are given in Table I. In addition, to ensure operation within the linear dynamic range of the VCSEL, the peak-to-peak amplitude (i.e., standard deviation) of the OFDM signal is set to V. This value is acquired by sweeping in an appropriate range according to the light-current-voltage (LIV) characteristics of the VCSEL and specifying the value that maximizes the link data rate by post-processing the measurement data in MATLAB.
[…]
@@
Fig. 5 shows the average BER performance with respect to the achievable data rate. It can be observed that the achievable data rate increases steadily at the cost of increasing the BER at the same time. It can also be observed that Config. I performs slightly better than Config. II. A possible explanation for this is Config. I sets the VCSEL’s DC bias at a lower operating point with slightly smaller voltage and current values. Therefore, the modulating signal experiences less nonlinear distortion effects, particularly for the signal swing above the DC operating point. In other words, in Config. I, the OFDM signal variations match the linear dynamic range of the VCSEL. As a result, the SNR is slightly improved especially in the lower frequency region up to GHz, as shown in Fig. 3, thereby improving the bit loading performance, as shown in Fig. 4. We consider a target BER of according to system-level requirements in LTE systems. In fact, in LTE systems, link adaptation aims to maintain a block error rate (BLER) of after decoding [14, Chapter 10, pp. 215–248]. This allows the system to tolerate a relatively high raw, uncoded BER on the order of to at the physical layer depending on the selected modulation and coding scheme (MCS) and the operating SNR so as to maximize spectral efficiency, relying on advanced FEC coding schemes such as Turbo codes or hybrid automatic repeat request (HARQ) mechanisms to enhance the BER performance [15]. The final post-FEC BER is expected to be at or lower after decoding with to overhead [16]. The BER performance can be further reduced to after decoding by using staircase codes with a overhead [16]. From Fig. 5, it can be observed that the achievable data rate exceeds Gb/s while . More precisely, the VCSEL-based optical wireless link achieves data rates of Gb/s and Gb/s at and under Config. I and Config. II, respectively. We note that both configurations achieve a data rate above Gb/s at a pre-FEC BER level of , which offers a improvement relative to the data rate achieved in [6] at the same BER level.

### v2

@@ Abstract
I Introduction
@@
Achieving data rates of tens of Gb/s using laser-based OWC for indoor applications has been actively investigated [4, 5, 6]. In [4], Lee et al. experimentally validated a dual‑wavelength wavelength division multiplexing (WDM) OWC system using a novel white‑light transmitter by integrating a nm gallium nitride (GaN) blue laser, an nm infrared (IR) laser, and a phosphor converter within a single surface‑mount device (SMD) package. The authors reported a combined data rate exceeding Gb/s over a m link at a target bit error ratio (BER) of adopted from long term evolution (LTE) systems, using two silicon (Si) positive-intrinsic-negative (PIN) photoreceivers with a dB bandwidth of GHz to detect the optical signals of the blue and IR laser diodes. In [5], Chen et al. implemented a ‑channel LiFi link using the same SMD laser technology based on a transmitter setup that integrates ten high‑brightness, high‑bandwidth LDs, including three GaN blue LDs operating at nm, nm, and nm, and seven IR LDs with wavelegths of nm to nm. To detect the optical signals, GHz Si PIN photoreceivers were used, except for the nm LD, for which a GHz indium gallium arsenide (InGaAs) photoreceiver was used due to its higher responsivity. The authors demonstrated aggregate data rates in excess of Gb/s over a few meter link distance by considering a BER threshold of .
[…]
@@ II-A VCSEL Characterization
We use an SM VCSEL operating at a center wavelength of nm with a low spectral width. This VCSEL was designed, manufactured and packaged by Vertically Integrated Systems (VIS) GmbH and was tested at the wafer level before being fully characterized for optical communications [13]. It is a multi-aperture (MA) VCSEL featuring a compact mini-array structure, comprising four oxide-confined VCSELs, each with an effective aperture diameter of about µm, which are arranged in a array configuration with a small pitch of µm [13]. These four VCSELs have a common pair of contacts such that they act as a single VCSEL. Consequently, the output power of the MA VCSEL is effectively improved by a factor of four while maintaining both the SM emission spectrum and the broadband modulation response of individual VCSELs in the mini-array structure [13]. The small-pitch VCSEL mini-array has a compact footprint for efficient light coupling into a µm core MMF [7]. Such a miniature structure is fabricated using oxidation through holes [7]. A die shot of the VCSEL mini-array chip is available in [13].
Fig. 1 shows the light-current-voltage (LIV) characteristics of the VCSEL measured at room temperature. The VCSEL has a maximum output power of mW, a forward threshold current of about mA, and a roll-over current of mA. In addition, it provides a linear input dynamic range with a slope efficiency of W/A for forward currents between mA and mA, which is suitable for DCO-OFDM modulation. Fig. 2 shows the optical spectrum of the VCSEL for different values of the direct current (DC) bias current , measured using a high-resolution optical spectrum analyzer (OSA). It can be observed that when increases from mA to mA, the peak wavelength shifts slightly from nm to nm. This spectral shift is mainly due to thermal effects, which cause the longitudinal resonance of the cavity to move to longer wavelengths [7]. For each case, the average wavelength and the root mean square (RMS) spectral width are also provided in Fig. 2. The RMS spectral width is a measure of the effective wavelength spread of the optical spectrum and is defined as the weighted standard deviation of the wavelength around its mean [14]:
where is the emitted power at wavelength . The value of directly determines the chromatic dispersion (CD) penalty and maximum reachable distance in MMF links [7]. According to Fig. 2, by increasing from mA to mA, increases from nm to nm. As a reference for comparison, the IEEE 802.3 standard for Ethernet over MMF links specifies a maximum RMS spectral width of nm for VCSELs operating at nm to nm to reach link distances up to m [15]. Fig. 3 shows the FF intensity profile of the VCSEL measured at different values of , which confirms that the VCSEL has a nearly Gaussian beam profile, with a full-width at half-maximum (FWHM) divergence angle ranging between and when increasing from mA to mA. The corresponding divergence half-angle, defined for the beam spot containing of the transmitted power, varies between and . Furthermore, the VCSEL has a -dB modulation bandwidth of GHz.
II-B End-to-End System Setup
@@
Fig. 4 shows the link configuration including the optical and optomechanical assemblies of the transmitter and receiver. Fig. 5 depicts the complete experimental setup. In this work, considering the low-pass and frequency-selective characteristics of the VCSEL-based OWC system, we use DCO-OFDM with adaptive subcarrier bit and power allocation to ensure high spectral efficiency. The modulation and demodulation of the OFDM signal are implemented based on -point inverse fast Fourier transform (IFFT) and FFT operations, respectively, where , using offline digital signal processing (DSP) in MATLAB running on a laptop computer. At the transmitter, the information bits are mapped to -ary quadrature amplitude modulation (QAM) symbols to be loaded across a total of data-carrying subcarriers, with the constellation size reaching up to , depending on the available signal-to-noise ratio (SNR). The DCO-OFDM frame is constructed by concatenating the block of the complex QAM symbols with its conjugated version according to a Hermitian symmetry. A cyclic prefix (CP) of length is then added and an IFFT is performed on the resulting frame to generate the time domain OFDM signal. This signal is further shaped using a root-raised cosine filter with a roll-off factor of 0.1. The pulse-shaped digital waveform is converted into a wideband analog signal via the Keysight M8195A arbitrary waveform generator (AWG), operating at a sampling frequency of GSa/s. According to the Nyquist-Shannon theorem, the bandwidth of the signal is related to the sampling frequency via . In practice, the signal bandwidth at the AWG output can be controlled using the number of samples per symbol (SPS) parameter, denoted by , based on . We choose to obtain GHz. The AWG output signal is combined with the required DC bias using a Mini Circuits ZX85-12G-S+ bias tee (BT) to drive the VCSEL. This BT has a GHz bandwidth with a minimum operating frequency of kHz, and an insertion loss (IL) of dB. The wideband OFDM signal is fed into the BT using a radio frequency (RF) coaxial cable with a bandwidth of GHz, a length of cm, and dB over the desired bandwidth. To collimate the light beam at the transmitter, a Thorlabs ACL5040U-B aspheric condenser lens is placed in front of the VCSEL. This lens has a diameter of mm with a clear aperture (CA) of mm, a focal length of mm, a numerical aperture (NA) of , and an anti-reflective coating (ARC) for – nm. Note that when the signal is directly modulated on the forward current of the MA VCSEL for intensity modulation (IM), although the output beams of the four apertures are overlapping while propagating in free space, there is no crosstalk between them, since they carry the same signal and their signal contributions are constructively added together upon direct detection (DD) at the receiver.
@@
At the receiver side, a Thorlabs C80FC-B achromatic fiber collimator (AFC) is employed to collect the incident beam from the transmitter located at a cm distance. This AFC has a CA of mm, an effective focal length of mm, an ARC for – nm matching the coating of the transmitter lens, and a Ferrule Connector/Physical Contact (FC/PC) fiber-optic output with a maximum fiber NA of . As shown in Fig. 4, the AFC is mounted on a three-axis kinematic mount with high-resolution angular adjusters along with an coordinates translation stage, providing six degrees of freedom (DOF) required for precise alignment of the focused beam spot on the fiber plane to ensure high coupling efficiency. The output port of the AFC is connected to the optical input of the fiber-coupled receiver using a graded-index OM4 MMF patch cable of length m with a core diameter of µm and FC/PC connectors. We use a Thorlabs RXM25BF fiber-optic photoreceiver, which is equipped with an InGaAs PIN PD followed by integrated transimpedance amplifier (TIA) and variable gain amplifier (VGA) stages. This receiver has a broadband responsivity in – nm, and a photodetection bandwidth of up to GHz with a minimum operating frequency of kHz. The built-in VGA helps with adjusting the transimpedance gain and the peak of the receiver frequency response to optimize the performance. There is also a digital screen on the photoreceiver to monitor the average detected photocurrent, which helps with performing the alignment using an unmodulated laser beam. The output analog signal from the photoreceiver is converted to a digital signal by the Keysight UXR0104B oscilloscope at a sampling frequency of GSa/s. The remaining DSP is performed in MATLAB. This includes frame synchronization, FFT, single-tap equalization, and QAM demapping. For each experiment, a total of OFDM frames consisting of pilot frames and data frames are sent through the OWC system to measure the BER performance. As shown in Fig. 4, the back-to-back reference signal provides a copy of the transmitted OFDM signal on the oscilloscope Channel 2 through a cable connection with the AWG, which can be used as the triggering channel to capture the waveform of the received OFDM signal.
Since the AFC is mounted by using an optomechanical assembly with six DOF, any misalignment including lateral, angular, or axial, has an impact on the light coupling into the MMF core. It is particularly sensitive to angular misalignments when adjusting the three tilt angles of the AFC. Even a tiny turn of one of the adjusting knobs on the kinematic mount can greatly shift the beam spot on the focal plane of the AFC, causing the coupling efficiency to drop dramatically, which makes the alignment process practically challenging. We choose a link distance of cm to alleviate the complexity of alignment and achieve reliable performance. However, the proposed setup can be systematically reconfigured to increase the link distance. Note that the end-to-end performance is not limited by fiber dispersion or attenuation due to the short length of the OM4 MMF cable. To elaborate, Appendix A provides the maximum bandwidth and the maximum reachable distance of MMF links under both CD and modal dispersion (MD). The fiber patch cable has an effective modal bandwidth (EMB) of MHzkm and a loss of dB/km at nm [16]. We adopt these values to present a conservative estimate of the performance at nm, since EMB increases and decreases with wavelength [17]. Based on (14), by substituting m, ps/nmkm [18], and nm from Fig. 2, we obtain THz. Compared to the signal bandwidth GHz, , indicating that CD and MD have practically no impact on the performance. Alternatively, based on (15), we find m, which leads to the same conclusion. In addition, the fiber attenuation for m is dB, which is quite negligible.
II-C Achievable Data Rate
@@
Hence, the achievable data rate in (2) is upper-bounded as:
| [V] | [mA] | [mW] | [µW] | |
| Config. I | ||||
| Config. II |
: Transmitted optical power.
@@ III Measurement Results
We present the measurement results for the performance of the VCSEL-based OWC system in terms of the received SNR, average BER and link data rate. The channel transfer function is estimated by sending pilot -QAM symbols of unit variance on all data-carrying subcarriers [5]. The estimated channel information is then used to determine the optimal average transmit power and constellation order of QAM symbols for each subcarrier based on the corresponding subcarrier SNR, thus adapting the transmission rate and maximizing the system spectral efficiency. In this work, the Hughes-Hartogs (HH) bit and power loading algorithm is applied [21], which is widely used in multi-carrier communication systems. The optimal DC operating point of the VCSEL is obtained through extensive measurements and two configurations are identified. These configurations along with the resulting average received optical powers, denoted by , are given in Table I. In addition, to ensure operation within the linear dynamic range of the VCSEL, the peak-to-peak amplitude (i.e., standard deviation) of the OFDM signal is set to V. This value is acquired by sweeping in an appropriate range according to the LIV characteristics of the VCSEL and specifying the value that maximizes the link data rate by post-processing the measurement data in MATLAB.
III-A SNR Frequency Response: Adaptive Bit and Power Loading
Fig. 6 shows the channel gain and the electrical SNR as a function of frequency. This channel gain represents the channel transfer function which is measured in the frequency domain between the IFFT input at the transmitter and the FFT output at the receiver. It can be observed that the channel gain exhibits a sharp decline after GHz. This is primarily caused by the hard cutoff frequency of the oscilloscope at about GHz. Although the VCSEL and the fiber-coupled receiver can reach a higher bandwidth (i.e., GHz and GHz, respectively), in this experiment, the oscilloscope limits the bandwidth of the end-to-end system. We note that since the BT has a bandwidth of GHz, it also causes a roll-off on the frequency response at high frequencies. However, the cutoff frequency at GHz is still dominant. Evidently, the resulting SNR values are above dB over the entire range of frequencies up to GHz. Thus, the effective modulation bandwidth that can be used for data transmission on OFDM subcarriers is GHz. It can be observed that for GHz, dB for GHz, and dB for GHz. Furthermore, there are oscillations in the SNR response especially at frequencies lower than GHz, due to the impedance mismatch between the VCSEL and the BT. In fact, while the BT has an output impedance of Ω, the input impedance of the VCSEL is significantly less than Ω (i.e., about Ω). Although the VCSEL chip is packaged on an standard RF-V connector and it is directly connected to the BT output, as shown in Fig. 4, its low input impedance leads to signal reflections at this critical connection, resulting in RF interference that degrades the SNR and hence oscillations appear in the SNR frequency response. The remaining cable connections along the RF signal path at the transmitter and receiver are carefully chosen to be as short as possible to minimize RF signal distortions.
@@ III-B BER vs Achievable Data Rate
Fig. 8 shows the average BER performance with respect to the achievable data rate. It can be observed that the achievable data rate increases steadily at the cost of increasing the BER at the same time. It can also be observed that Config. I performs slightly better than Config. II. A possible explanation for this is Config. I sets the VCSEL’s DC bias at a lower operating point with slightly smaller voltage and current values. Therefore, the modulating signal experiences less nonlinear distortion effects, particularly for the signal swing above the DC operating point. In other words, in Config. I, the OFDM signal variations match the linear dynamic range of the VCSEL. As a result, the SNR is slightly improved especially in the lower frequency region up to GHz, as shown in Fig. 6, thereby improving the bit loading performance, as shown in Fig. 7. We consider a target BER of according to system-level requirements in LTE systems. In fact, in LTE systems, link adaptation aims to maintain a block error rate (BLER) of after decoding [22, Chapter 10, pp. 215–248]. This allows the system to tolerate a relatively high raw, uncoded BER on the order of to at the physical layer depending on the selected modulation and coding scheme (MCS) and the operating SNR so as to maximize spectral efficiency, relying on advanced FEC coding schemes such as Turbo codes or hybrid automatic repeat request (HARQ) mechanisms to enhance the BER performance [23]. The final post-FEC BER is expected to be at or lower after decoding with to overhead [24]. The BER performance can be further reduced to after decoding by using staircase codes with a overhead [24]. From Fig. 8, it can be observed that the achievable data rate exceeds Gb/s while . More precisely, the VCSEL-based optical wireless link achieves data rates of Gb/s and Gb/s at and under Config. I and Config. II, respectively. We note that both configurations achieve a data rate above Gb/s at a pre-FEC BER level of , which offers a improvement relative to the data rate reported in [6] at the same BER level.
III-C Achieving Gb/s Through SNR Extrapolation
Looking at Fig. 6, a natural question arises as to how much performance enhancement can be achieved if the oscilloscope bandwidth constraint is eliminated? To address this question, we estimate the trend of the SNR in the frequency domain by approximating the measured SNR frequency response based on piecewise linear (PWL) regression. For each configuration, the approximated SNR is characterized by two frequencies of and such that , where GHz indicates the frequency at which the SNR drops to dB, as shown in Fig. 9. A moving average filter with a window size of is also employed to smooth out the ripples in the measured SNR results due to RF reflections. The resulting curve manifests the SNR trend and is used as a guide for line fitting in each frequency band. The computed values of and for both configurations are listed in Table II. The approximated SNR is then extrapolated by removing the hard limit of the oscilloscope at and letting the SNR to follow the same decreasing trend as in the frequency band until it crosses dB at , as shown in Fig. 9. This leads to GHz for Config. I, and GHz for Config. II. The rationale behind this extrapolation lies in the fact that the frequency response of SM MA VCSELs exhibits a monotonically decreasing behavior with a consistent roll-off at high frequencies [25]. Thus, when there is no other factor limiting the end-to-end modulation bandwidth, the system frequency response follows a similar decreasing trend.
| [GHz] | [GHz] | [GHz] | |
| Config. I | |||
| Config. II |
For an additive white Gaussian noise (AWGN) channel with an SNR frequency response of , the achievable data rate can be upper-bounded as [26]:
where is given by (4). When evaluating the upper bound in (6) for the approximated and extrapolated SNR, we have:
where is replaced by for the approximated SNR, and by for the extrapolated SNR. Fig. 10 shows the BER performance with respect to the achievable data rate for both cases based on the upper bound in (7). As observed from Fig. 10, under the approximated SNR, the upper bound of the data rate gets tighter for higher BER values, indicating that it can be safely used to estimate the maximum achievable data rate. With the extrapolated SNR, the achievable data rate is remarkably enhanced, reaching a maximum value of Gb/s and Gb/s at a BER of and for Config. I and Config. II, respectively. This confirms that the system can achieve data rates Gb/s.